# Eaman Jahani

### v4

@@ F.1 Hypotheses and Results
    Mean prompt length (words): DALL-E 3 Revised has significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
    Prompt embedding variance: DALL-E 3 Revised also has significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
    Cosine similarity with target image, Z-Score: DALL-E 3 Revised has significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
@@
    Mean percentile rank of CosineSim: DALL-E 3 (verbatim) has significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7352, $p=0.0009$).
    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7888, $p=0.0141$).
    Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7961, $p=0.0159$).
    DreamSim with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8195, $p=0.0377$).
    Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) has significantly more variance (ratio 1.2303, $p=0.0301$).
@@
    Mean prompt length (words): DALL-E 3 (verbatim) has significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
    Prompt embedding variance: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
    Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7501, $p=0.0015$).
    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7904, $p=0.0141$).
    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8005, $p=0.0159$).
@@
QTE Highlighted Results: Through our QTE plots, we see clearest evidence of dispersion being reduced for:
  • Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALLE-3 Revised vs. DALLE-2 comparison.
And we see the clearest evidence of dispersion increasing for:
  • Mean prompt length for all three pairwise model comparisons (DALLE-3 Verbatim vs. DALLE-2, DALLE-3 Revised vs. DALLE-2, and DALLE-3 Verbatim vs. DALLE-3 Revised)
  • Raw Dreamsim performance for DALLE-3 Verbatim vs. DALLE-2 and DALLE-3 Revised vs. DALLE-2 comparisons.
@@
Analysis approach:
In order to test for differences in prompting performance improvement across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata. The two samples represent the best and initial scores. We make these comparisons overall and within strata.
@@
    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
@@
    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
@@
Analysis approach:
To evaluate this hypothesis, we pre-registered and performed two analyses. First, we compared the extent to which different images can be replicated using GPT-4V—a multimodal generative AI model that can take in both text and images and output text. For each of the 15 target images, we prompted GPT-4V to “Write a DALL-E {2, 3} prompt to recreate this image verbatim as closely and as detailed as possible.” We do this once for DALL-E 2 and once for DALL-E 3, generating two “AI prompts” per target image. We then provided these AI prompts to DALL-E 2, DALL-E 3 (Verbatim), and DALL-E 3 (Revised), respectively, 20 times each, generating 60 replicated images per target image (the DALL-E 3 AI prompt was sent to both DALL-E 3 (Verbatim) and (revised)). For each image, we measure the cosine similarity of the CLIP embedding vectors to those of the relevant target image. We then average these 60 similarity measures by target image, resulting in a mean similarity score that represents GPT-4V’s ability to generate prompts that recreate each target image. Second, we simply measured the similarity of every participant-generated image to that of the relevant target image, and averaged over these participant-generated similarities.
@@ Results:
  • CLIP embedding cosine similarity: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
@@
    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for GPT-4V to replicate) was $CoSim=0.944$ for Business Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $CoSim=0.734$ for Photography Image #4.
    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
@@
    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, and the lowest score (the hardest image for participants to replicate was $CoSim=0.669$ for Business Image #4.
  • DreamSim: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
@@
    The maximum 1-DreamSim score in the above list (the easiest image for GPT-4V to replicate) was $\tilde{D}=0.75$ for Design Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $\tilde{D}=0.40$ for Photography Image #4.
    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
@@
    The maximum 1-DreamSim score in the above list (the easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, and the lowest score (the hardest image for participants to replicate was $\tilde{D}=0.356$ for Photography Image #4.
@@
In our pre-registration, we also mention a number of exploratory analyses that are not described with the same level of detail. Multiple of these exploratory analyses appear in our main text. These pre-registered exploratory analyses are copied verbatim below:
“We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate. We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
@@
F.2 Deviations From Pre-registration
Here, we also report any deviations from our pre-registered analysis. By and large, these deviations occurred because certain aspects of our pre-registration were not appropriate from a statistical analysis perspective, or were infeasible.
@@ F.2 Deviations From Pre-registration
  • In our pre-registration, we had anticipated using t-tests and Mann-Whitney U tests for many hypotheses. However, this turned out to be impossible for many variables since most of the demographic traits have multiple categories. Thus, we applied ANOVA and Kruskal-Wallis tests instead.
@@
  • In our pre-registration, we were not explicit enough on how to compute Z-score measures of performance. As described in section C.2, Z-scores are almost always computed within image-attempt to adjust for variations across target images or attempts. The only exception is with any analysis that compares performance across attempts where Z-scores need to computed within target images. The pre-registration anticipated that Z-scores for any analysis would be computed only within target images.
  • In our pre-registration, we had forgotten to include Education and GenAI outlook in our list of demographics with respect to which we would measure task performance heterogeneity. We decided to include these important variables in our analysis, despite the accidental omission.
@@
  • In addition to our pre-registered prompt exclusion criteria, we removed additional prompts that did not appear to be “good-faith efforts” to complete our task based on the text of those prompts (see Section B.4 for more details). We find that our results are the robustness to the inclusion of prompts that were removed by this procedure.
@@
  • Because they were easier to implement, we ran t-tests rather than z-tests to conduct our tests of H5. Because the two tests are asymptotically equivalent, we do not believe this will make a difference in our analysis.
  • When conducting our analyses to test H4, we observed that the distribution of the Z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values. To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute Z-scores greater than 3, which constituted 2.65% of the user-attempt observations. For robustness, we also conducted the stratification analysis including these outliers and found almost identical results, with the exception that the difference in performance between the top bracket and middle or bottom brackets was in some cases only marginally significant.
@@
  • In our pre-registration, we had included a rescaled version of our main performance metrics described in B.5.1 based on GPT-4V for robustness check. This outcome variable would have been rescaled by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image. Given that the quality of the prompt from GPT-4V model is a constant for each target image, this would have amounted to subtracting a constant from the unscaled CLIP cosine similarity or dream sim scores. Furthermore, since all our analysis involving performance as the dependent variable adjust for the target image, either through post-stratification or as a covariate in a linear model, the results from using GPT-4V scaling would have been identical to the the unscaled measures of performance. For this reason, we have not included the robustness checks involving GPT-4V for our pre-registered hypothesis.

### v5

@@ G.1 Hypotheses and Results
    • Mean prompt length (words): DALL-E 3 Revised showed significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
    • Prompt embedding variance: DALL-E 3 Revised also showed significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
    • Cosine similarity with target image, Z-Score: DALL-E 3 Revised showed significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
@@
    • Mean percentile rank of CosineSim: DALL-E 3 (verbatim) showed significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
    • Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also showed significantly less variance (ratio 0.7352, $p=0.0009$).
    • Mean percentile rank of DreamSim: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7888, $p=0.0141$).
    • Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7961, $p=0.0159$).
    • DreamSim with target image, Z-Score: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.8195, $p=0.0377$).
    • Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) showed significantly more variance (ratio 1.2303, $p=0.0301$).
@@
    • Mean prompt length (words): DALL-E 3 (verbatim) showed significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
    • Prompt embedding variance: DALL-E 3 (verbatim) also showed significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
    • Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7501, $p=0.0015$).
    • Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.7904, $p=0.0141$).
    • Mean percentile rank of DreamSim: DALL-E 3 (verbatim) showed significantly less variance (ratio 0.8005, $p=0.0159$).
@@
QTE Highlighted Results: Through our quantile treatment effects analysis and visual inspection of QTE plots, we found:
  • Evidence of dispersion reduction: Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALL-E 3 Revised vs. DALL-E 2 comparison showed patterns consistent with inequality reduction, with larger positive effects at lower quantiles.
  • Evidence of dispersion increase: Mean prompt length for all three pairwise model comparisons (DALL-E 3 Verbatim vs. DALL-E 2, DALL-E 3 Revised vs. DALL-E 2, and DALL-E 3 Verbatim vs. DALL-E 3 Revised) showed patterns consistent with increased inequality, with larger positive effects at higher quantiles. Similar patterns emerged for raw DreamSim performance in both DALL-E 3 Verbatim vs. DALL-E 2 and DALL-E 3 Revised vs. DALL-E 2 comparisons.
These findings present a nuanced picture of how model capacity affects variability in performance and prompting behaviors. For prompt length and embedding variance, the more advanced models (both DALL-E 3 variants) showed significantly less variance than DALL-E 2, suggesting more consistent prompting behaviors. For performance measures, the results were mixed: DALL-E 3 (Verbatim) generally showed less variance than DALL-E 3 (Revised), indicating more consistent performance across participants, while DALL-E 3 (Revised) showed higher variance in performance than DALL-E 2.
The QTE analysis suggests that the relationship between model capacity and performance inequality is complex. For some measures (e.g., Z-score of cosine similarity), more advanced models appeared to reduce inequality by disproportionately benefiting lower-performing users. For other measures (e.g., prompt length and raw DreamSim scores), more advanced models appeared to increase inequality by disproportionately benefiting higher-performing users.
@@
Analysis approach: To evaluate how performance improves across successive attempts and whether this improvement varies by model, we conducted stratified two-sample tests comparing participants’ initial scores with their best scores. These comparisons were performed both within each treatment arm (DALL-E 2, DALL-E 3 Verbatim, and DALL-E 3 Revised) and overall across all participants. This approach allowed us to determine both whether participants generally improved with practice and whether the rate of improvement differed across models with varying capabilities.
Results: Our analyses revealed consistent performance improvements across all treatment arms and outcome measures:
@@
    • No rescaling: We found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample. This indicates that participants were able to improve their performance through iterative attempts regardless of which model they were using.
    • Z-score rescaling: Results mirrored those of the unscaled analysis, with statistically significant improvements from initial to best performance observed both within each treatment arm and overall.
    • Percentile rank rescaling: Consistent with other scaling approaches, we observed statistically significant improvements from initial to best performance both within each treatment arm and overall.
@@
    • No rescaling: DreamSim results aligned with CLIP similarity findings, showing statistically significant improvements from initial to best performance both within each treatment arm and overall.
    • Z-score rescaling: We again found statistically significant improvements from initial to best performance both within each treatment arm and in the overall sample.
    • Percentile rank rescaling: Results were consistent with other analyses, showing statistically significant improvements from initial to best performance across all conditions.
These findings demonstrate that participants consistently improved their performance through iterative attempts across all model conditions and performance measures. Users were able to learn from feedback and refine their prompts to achieve better results over time, regardless of which model they were using. This learning effect was robust across different outcome measures and scaling approaches, highlighting the importance of iteration and feedback in developing effective prompting strategies.
Although our original hypothesis also posited that the rate of improvement might vary as a function of model capacity, we did not find consistent evidence for differences in improvement rates across models. This suggests that while more capable models yield better absolute performance (as demonstrated in H5), the relative improvement from first to best attempt was similar across model conditions. This finding highlights that the benefits of iterative refinement apply broadly across different model capabilities.
@@
Analysis approach: To evaluate this hypothesis, we employed two complementary approaches. First, we used GPT-4V (a multimodal generative AI model capable of processing both text and images) to generate optimal prompts for each target image. For each of our 15 target images, we instructed GPT-4V to ”Write a DALL-E 2, 3 prompt to recreate this image verbatim as closely and as detailed as possible.” We generated two such ”AI prompts” per target image—one optimized for DALL-E 2 and one for DALL-E 3. We then submitted these AI-generated prompts to the respective models 20 times each, yielding 60 replicated images per target (as the DALL-E 3 prompts were sent to both DALL-E 3 variants). We measured the cosine similarity between CLIP embeddings of these generated images and their target images, then averaged these similarity scores to quantify GPT-4V’s ability to generate effective replication prompts for each target.
Second, we analyzed human performance by measuring the similarity between all participant-generated images and their corresponding target images, averaging these similarities to determine which images were easier or harder for participants to replicate. This two-pronged approach allowed us to assess image difficulty from both AI and human perspectives.
@@ Results:
  • CLIP embedding cosine similarity: When ranking target images by GPT-4V’s ability to generate effective prompts, we found substantial variation across images:
@@
    The highest average cosine similarity (easiest image for GPT-4V to replicate) was $CoSim=0.944$ for Business Image #3, while the lowest score (hardest image to replicate) was $CoSim=0.734$ for Photography Image #4.
    When ranking target images by participants’ ability to generate effective prompts, we observed:
@@
    The highest average cosine similarity (easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, while the lowest score (hardest image to replicate) was $CoSim=0.669$ for Business Image #4.
  • DreamSim: When using DreamSim as our similarity metric (inverted as 1-DreamSim to create a similarity rather than distance measure), GPT-4V’s prompt generation performance ranked as follows:
@@
    The highest 1-DreamSim score (easiest image for GPT-4V to replicate) was $\tilde{D}=0.75$ for Design Image #3, while the lowest score (hardest image to replicate) was $\tilde{D}=0.40$ for Photography Image #4.
    When ranking target images by participants’ replication ability using DreamSim:
@@
    The highest 1-DreamSim score (easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, while the lowest score (hardest image to replicate) was $\tilde{D}=0.356$ for Photography Image #4.
These results clearly demonstrate substantial variation in image replicability across our target set. Business-related images were generally easier to replicate, while photography images tended to be more challenging. While the exact rankings varied somewhat between GPT-4V and human participants, and between similarity metrics, the overall pattern of relative difficulty remained fairly consistent. This variation in difficulty confirms the importance of our stratified analysis approach and underscores the need to consider image-specific characteristics when evaluating generative AI performance.
@@
Our pre-registration outlined several exploratory analyses that we planned to conduct beyond our primary hypotheses. Several of these exploratory analyses appear in our main text, particularly the replay analysis that decomposed performance improvements into model and prompting effects. The pre-registered exploratory analyses were described as follows:
”We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate. We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
@@
As noted in the main text, the third exploratory analysis—submitting prompts from one model condition to another model—became central to our investigation of model versus prompting effects. This approach allowed us to isolate how much improvement came from the model’s enhanced capabilities versus users adapting their prompting strategies to take advantage of those capabilities.
G.2 Deviations From Pre-registration
We report below all deviations from our pre-registered analysis plan. These deviations primarily resulted from statistical or methodological considerations that became apparent during data analysis, rather than from substantive changes to our research questions or hypotheses.
@@ G.2 Deviations From Pre-registration
  • Statistical tests: Our pre-registration specified t-tests and Mann-Whitney U tests for many hypotheses. However, this approach proved inappropriate for variables with multiple categories (which characterized most demographic traits). We therefore employed ANOVA and Kruskal-Wallis tests instead, which provide equivalent information for multi-category variables.
@@
  • Z-score computation: Our pre-registration was insufficiently precise about computing Z-score performance measures. As detailed in Section D.2, we calculated Z-scores within image-attempt pairs to adjust for variations across target images and attempts, except when comparing performance across attempts, where Z-scores were computed within target images only. The pre-registration had anticipated computing Z-scores only within target images for all analyses.
  • Additional demographic variables: We inadvertently omitted Education and Generative AI outlook from our pre-registered list of demographics for measuring task performance heterogeneity. Given their theoretical importance, we included these variables in our analyses despite this accidental omission.
@@
  • Additional prompt exclusion criteria: Beyond our pre-registered exclusion criteria, we removed additional prompts that did not appear to constitute ”good-faith efforts” based on their text content (see Section C.4 for details). Robustness checks confirmed that our results remain consistent when including these prompts.
@@
  • T-tests vs. Z-tests: For testing H5, we used t-tests rather than z-tests because they were easier to implement. As these tests are asymptotically equivalent, this change does not meaningfully affect our results.
  • Z-score outlier handling: When testing H4, we observed that Z-scores did not follow a normal distribution and included extreme values (ranging from -21.9 to 7.1). To prevent these outliers from disproportionately influencing our linear models, we excluded observations with absolute Z-scores greater than 3 (2.65
@@
  • GPT-4V rescaling omission: Our pre-registration included a rescaled version of performance metrics based on GPT-4V prompt quality. However, this rescaling would have amounted to subtracting a constant from each unscaled score, and since all our analyses adjusted for target image (either through post-stratification or as a covariate), results would have been identical to those using unscaled measures. We therefore omitted these redundant robustness checks.
G.3 OSF Pre-registration
A complete copy of the pre-registration deposited on OSF on December 11, 2023 will be included beginning on the next page in the final manuscript, but has been removed here to maintain anonymity.